### v1

Abstract
The deployment of autonomous agents in environments involving human interaction has increasingly raised security concerns. Consequently, understanding the circumstances behind an event becomes critical, requiring the development of capabilities to justify their behaviors to non-expert users. Such explanations are essential in enhancing trustworthiness and safety, acting as a preventive measure against failures, errors, and misunderstandings. Additionally, they contribute to improving communication, bridging the gap between the agent and the user, thereby improving the effectiveness of their interactions. This work presents an accountability and explainability architecture implemented for ROS-based mobile robots. The proposed solution consists of two main components. Firstly, a black box-like element to provide accountability, featuring anti-tampering properties achieved through blockchain technology. Secondly, a component in charge of generating natural language explanations by harnessing the capabilities of Large Language Models (LLMs) over the data contained within the previously mentioned black box. The study evaluates the performance of our solution in three different scenarios, each involving autonomous agent navigation functionalities. This evaluation includes a thorough examination of accountability and explainability metrics, demonstrating the effectiveness of our approach in using accountable data from robot actions to obtain coherent, accurate and understandable explanations, even when facing challenges inherent in the use of autonomous agents in real-world scenarios.
keywords:
[…]
{highlights}
Security improvements for autonomous agents in public spaces.
Review of blockchain trends for immutable logs, EDRs, XARs, and LLMs.
Accountability and Explainability architecture for autonomous agents.
Comprehensive assessment across diverse navigation scenarios and metrics.
Future research directions for accountable and explainable autonomous agents.
1 Introduction
The rising presence of mobile robots alongside humans has increased growing security and safety concerns. One of the main challenges in this context is the ability of robots to justify their actions and decisions, particularly in safety-critical scenarios (Sado et al., 2023). Therefore, it is essential to ensure that the events performed when developing an action or behavior are reliable and transparent (Borgo et al., 2018). Equally important is the capacity to render such explanations comprehensibly to non-expert users, offering appropriate, effective, and natural answers to their queries (Rosenfeld and Richardson, 2019; Cruz et al., 2023).
The most substantial evidence in post-event analysis relates to the logging and auditing capabilities, which are essential to mitigate technical issues and provide forensic evidence to detect any data tampering or unauthorized leakage. Therefore, a trustworthy and secure log system is needed when generating a complete trace of a robot’s behavior (Al-Slais and Ali, 2023). The system must ensure the integrity, confidentiality, and availability of the agent and the data. Additionally, it should possess the capability of incident detection, facilitating investigations, and ensuring compliance, ultimately improving the overall security of the robot.
Regarding log systems vulnerabilities, covering tracks has become a common practice for attackers, erasing any traces that could lead to their identification or assisting in the analysis of their actions. As a result, it is necessary to strengthen system properties against unauthorized modifications (Ahmad et al., 2019). This fact not only includes safeguarding the integrity of log data but also implementing measures to detect and counteract attempts at tampering. Accountability is equally required, enabling detection and isolation of faulty behaviors and their origin. It must not only provide tamper resistance, but also to offer verifiability for checking the integrity of log entries against tampering. Moreover, it should guarantee search accessibility, enabling access to specific records whose integrity can be verified (Chiu et al., 2021).
Integrating accountability solutions into robotic systems often demands a trade-off with system performance. Specifically, features such as computational power and networking bandwidth can be negatively affected. The inclusion of anti-tampering techniques, such as full disk and communication encryption, may be restrictive due to resource limitations (Fernández-Becerra et al., 2022). Additionally, high-cost tamper proof storage devices may prove economically unviable, especially when dealing with continuously generated log data.
Therefore, it becomes necessary to establish an efficient and verifiable proof of integrity to guarantee the reproducibility and identification of the events that could arise during an autonomous agent action. For this evidence to be significant, it should belong to a persistent chain of custody, totally reliable and verifiable, main requirements for an auditable and transparent logging system (Putz et al., 2019; Hajer et al., 2024).
Blockchain solutions emerge as a mechanism to ensure data integrity, protection against unauthorized changes, and traceability. The consensus protocols established by all blockchain network members enable the detection and rejection of any destruction, manipulation, or non-permitted editing of data through illicit channels. The nature of this technology helps to avoid tampering risk, safeguarding data (Iqbal and Matulevičius, 2019) or even .
Nevertheless, the potential obtained evidence poses significant challenges for analysis due to its diverse and varied nature, terminology, and the presence of substantial volumes of similar messages unrelated to anomalous events. These facts may result in limited usefulness and misalignment with the expectations of non-expert users (Shehu and Harper, 2022). Regarding autonomous agents, the raw log messages produced by these systems, are usually noisy and possess semi-structured format. In addition to this, it is necessary to deal with specific features such as large quantities of domain-specific language messages (Ott et al., 2021).
In this particular context, explaining the reasons that make a robot carry out a specific and unexpected action is essential for diagnosing failures in goal achievement. Besides, making these explanations accessible to humans is equally important. This requirement has led to the development of eXplainable Autonomous Robots (XAR), which aims to provide mechanisms for presenting clear and understandable explanations for the actions and decisions of autonomous agents. XAR enhances the transparency and interpretability of these agents’ decision-making processes, often incorporating techniques like natural language explanations to improve clarity (Anjomshoae et al., 2019; Langley et al., 2017).
Large Language Models (LLMs) have gained significant attention due to their exceptional performance in diverse natural language processing tasks, including translation, coding or text summarization (Kocmi and Federmann, 2023; Zhang et al., 2024; Sarsa et al., 2022). One approach to improve the results of large language models for a specific domain is through the use of retrieval-based methods. By retrieving relevant information from a large corpus of text, these models can provide explanations grounded in existing knowledge (Ren et al., 2023).
This work presents an accountability and explainability architecture implemented for ROS-based mobile robots. This solution included two main components. Firstly, a black box-like component to provide accountability that includes anti-tampering features by distributing integrity proofs from the events recorded by the black box component. The second component generates natural language explanations from data in the black box, including the use of interactive natural language conversations as a means to provide comprehensive explainability for the recorded events. Both components are connected through an additional component responsible for processing raw data collected by the first one, eliminating non-essential information and building connections and relations between data. These functionalities release software developers from the need to make calls to the log API to build explanations for the code generated and the behaviors developed by the autonomous agent. Additionally, we have conducted a thorough evaluation of our solution, including its impact on system performance and the correctness, accuracy, helpfulness and coherence of the generated natural language explanations.
The remainder of this paper is organized as follows: Section 2 provides an overview of previous works and findings related to the study, focusing on the enhancement of immutable data recorders and explainability solutions for robots. Section 3 delves into the system architecture, offering a detailed description of the designed approaches. Section 4 details the technical elements of the designed system attending to the proposed architecture. Section 5 discusses the experimental process, the scenarios and the variables employed for assessment. Section 6 outlines the obtained results, particularly addressing accountability and explainability metrics. Finally, Section 7 summarises the obtained conclusions.
2 Related works
In this section, we explore key works related to various domains relevant to our approach. It includes progress in understanding how blockchain ensures the immutability of logs, research on the application of event data recorders in robotics, and significant contributions Explainable Autonomous Robots (XAR) field. Lastly, we examine the evolution of Large Language Models (LLMs) to provide natural language explanations.
2.1 Blockchain for immutable logs
Blockchain is a decentralized solution that eliminates the necessity for trust in a central authority through its immutable and distributed ledger, composed of timestamped transaction blocks. These blocks are linked through hashing and storage of previous block content, ensuring the integrity of transactions. Any attempt to alter a transaction in one block requires the modification of all subsequent blocks, resulting in a computationally expensive process. As a result, data stored in the blockchain is immutable (Rajasekaran et al., 2022; Guo and Yu, 2022).
There are two primary types of blockchains based on access controls: permissionless blockchains allow open access and public transaction visibility, while permissioned blockchains limit participation and transaction visibility. Nevertheless, closed networks may pose security risks, as authorized users could recalculate subsequent blocks in the event of a block modification (Solat et al., 2021). While permissionless blockchains have traditionally related to scalability and throughput challenges, approaches like Ethereum have transitioned to Proof of Stake (PoS), a more energy-efficient, secure, accessible and scalable consensus protocol (Nguyen et al., 2019; Sanka and Cheung, 2021).
Blockchain technology ensures log integrity and auditability, offering unique advantages over alternatives such as distributed storage systems or IPFS (InterPlanetary File System), which may require central authorities for storage and verification (Omar et al., 2024). Blockchain offers consensus and immutability, coupled with the capability of integrating logic mechanisms such as smart contracts (SCs). SCs contain functions that either write values onto the blockchain, thus altering its state, or read data from it (Mohanta et al., 2018). These features make blockchain the preferred choice for applications where data integrity and trust are required (Nyaletey et al., 2019).
Different secure logging systems rely on permissioned blockchain frameworks for log immutability (Shekhtman and Waisbard, 2019), introducing accountability even in embedded systems (Chiu et al., 2021). To get a balance between throughput and cost-effectiveness, some works combine both categories (Zhao et al., 2023), developing approaches that include a permissioned blockchain to ensure per-entry immutability and public non-repudiation solution (Pourmajidi et al., 2023). Nevertheless, inherent features of permissionless blockchain such as transparency and trust, have also boosted the development of secure logging systems (Ge et al., 2019).
2.2 Robotic Event Data Recorders
Event Data Recorders (EDRs), often referred to as black boxes, have traditionally played a crucial role in in capturing and analyzing events within robotics systems. These recorders support the collection and storage of event data, helping to comprehend system behavior, detecting anomalies, and ensuring safety (Winfield et al., 2021, 2022). Therefore, these solutions are needed in forensic investigations, ethical assessments and transparency regarding autonomous agents (Taurer et al., 2018).
The integration of black box modules to enhance robots’ introspection capabilities contributes to improving system transparency by attributing responsibilities (Fernández-Becerra et al., 2022). These solutions typically include functionalities such as hashing, signing, and encryption (Schmidt et al., 2022). However, to ensure the integrity and authenticity of recorded events, anti-tamper features must also be in place (Srinivas Aditya et al., 2021).
Using blockchain technology in event data recorders can guarantee the integrity of a transparent and auditable record of events. These features enhance accountability and trust in the recorded data, as it becomes difficult for any party to manipulate or tamper with the recorded events without detection (Salah et al., 2019). Additionally, blockchain involves decentralized and distributed storage of event data, eliminating the necessity for central authorities or intermediaries and reducing the risk of data loss.
Further research explores the application of blockchain technology to improve robotic capabilities, particularly event data recording and management. The Black Block Recorder (BBR) system proposed in (White et al., 2019), employs permissioned blockchain technology for immutable black box logging in robotic applications. The BBR system combines Rosbag2 and distributed ledger technologies (DLTs) to ensure data integrity and immutability. The potential applications of blockchain in robot swarms, including secure communication, data logging, and consensus agreement are discussed in (Strobel et al., 2020). In (Lopes et al., 2021), a time-segmented consortium blockchain for robotic event registration is proposed, organizing and storing event records efficiently to allow for retrieval and analysis of specific events within specific time frames. Authors in (Zhang et al., 2022) present a tool to bridge the Robot Operating System (ROS) and the Ethereum blockchain, evaluating encryption efficiency and stability, transactions response time and packet loss rate. Nevertheless, these works lack capabilities to save integrity proofs in the blockchain at selected intervals, to enhance granularity, or including service definitions or smart contracts to improve the verifiability process.
2.3 Explainable Autonomous Robots
The need for explainability in autonomous robots arises from increasing complexity and autonomy of these systems, making it challenging for users to comprehend their behavior. Explainable Autonomous Robots (XARs) are designed to provide explanations for their actions and decisions to non-expert users. Such explanations enrich aspects like security and the trust level, preventing failures, errors, or conflicts in achieving objectives or undesired interactions (Han et al., 2019).
Despite the growing demand for explainability in autonomous agents, their use reveals a lack of practical and effective usability or interpretability in real environments (Abdul et al., 2018). Therefore, it is a problem that requires special attention, being necessary to explain the performed decisions. This explanation should include alternatives, choices made, expectations of each option, decision criteria, and any unexpected events that may have led to the modification of the initially planned behavior (Langley et al., 2017; Sadeghi et al., 2021).
As described in (Sakai and Nagai, 2022), an explanation regarding XAR, is an "abductive reasoning that bridges the gap between the individual explaining, which in this case may be a robot, and the person to whom they are explaining". Self-explanatory capabilities in social robots have been provided through the development of architectures and frameworks (Stange et al., 2022; Han et al., 2021).
Nevertheless, the essence of explainability regarding XAR lies in effective communication (Sakai and Nagai, 2022). Progress in natural language processing (NLP) technology through large language models (LLMs), based on the estimation of human statements and the relevance and position of the words contained in them, could improve the efficiency of human-robot interactions. This enhancement can create a sense that these systems truly comprehend and respond to individuals meaningfully, significantly benefiting their usefulness and applicability (Sado et al., 2023).
2.4 Large Language Models
Large Language Models (LLMs) have recently played a significant role in natural language processing, thanks to replacing recurrent neural networks with self-attention mechanisms (Vaswani et al., 2017). The design and implementation of the first Transformers models constitute the foundation of most LLM architectures. Trained on massive text data, LLMs include strong capabilities in natural language understanding and complex task solving (Zhao et al., 2023).
However, LLMs face two primary challenges. First, they tend to produce a significant amount of hallucinations due to their reliance on internal knowledge (weights) (Andriopoulos and Pouwelse, 2023). Second, their large parameter sizes and the associated high update costs, make traditional pre-training and fine-tuning methods unfeasible. As a solution, Retrieval-Augmented Generation (RAG) methods emerge as a solution for LLMs to effectively interact with the external world (Zhao et al., 2023). This strategy is particularly notable in question-answering systems, which represent the most widely adopted use cases for harnessing explanatory and reasoning capabilities of large language models. In this domain, RAG-based approaches are emerging as the standard for developing solutions that seamlessly integrate the advantages of a knowledge base (utilizing a vector store) and generative models.
LLMs display emergent abilities, such as in-context learning (ICL), where LLMs make predictions based on contexts augmented with a few examples, helping the model to learn from analogy (Wei et al., 2022). This paradigm simplifies inclusion of human knowledge into LLMs by modifying the prompt templates (Dong et al., 2022). ICL does not require altering the model’s parameters or training the model on a specific dataset; instead, it relies on guiding the LLM through the prompting interface to align its results with human values and preferences.
Nevertheless, evaluating large language models (LLMs) and the impact of previously described approaches, poses several challenges due to their complexity and the nature of their outputs. These difficulties can be attributed to several factors, including the lack of suitable evaluation metrics, biases in generated text, and the cost of human assessment (Chang et al., 2024). Nevertheless, recently, employing LLMs as evaluators for LLMs-based solutions has been proven to be nearly as effective as human judges (Zheng et al., 2023).
Previous works have demonstrated the potential of using LLMs to generate natural language descriptions of logging systems (Setianto et al., 2022) or robots’ behaviors (González-Santamarta et al., 2023), indicating room for improvements to deliver more accurate and complete interpretations. Nevertheless, to our knowledge, no previous research has investigated the capabilities of LLMs in generating natural language explanations within the context of an accountability approach for record interpretation.
3 System architecture
Our approach comprises to main components: a tamper-proof black box-like device, and a module responsible for generating natural language explanations for the data stored by the former. The first component offers services for building, storing, and verifying integrity proofs through smart contract specifications. The explainability component provides natural language explanations by implementing Retrieval-Augmented Generation (RAG), thereby enhancing the obtained results using the data recorded by the first component as an external source to enrich the context in the answer formulation process. Both components are integrated through the existence of two main asynchronous tasks focused on cleaning and processing the raw data generated by the black box-like device to increase the effectiveness of the explainability engine.
Figure 1 depicts a generic architecture of our proposal, highlighting in blue color the functionalities related to accountability, and in green those related to explainability. Below we provide a deeper description of the approaches designed and followed in order to develop and integrate the above mentioned components.
Figure 1: System architecture.
3.1 Black Box Recorder
Our first component focuses on recording accountable information that captures essential messages generated during the navigation task of an autonomous agent. The selected data will guarantee the reproducibility of the mentioned task, including details like the robot’s poses, the employed map, odometry data, linear and angular velocity, and laser scan readings.
The immutability of the previously described information is achieved through blockchain. Blockchain technology usually faces limitations regarding the size of the data that can be written into a single transaction. Efficiently storing autonomous agent messages becomes challenging due to these constraints. Our proposed solution involves selectively storing the hash of messages at intervals, providing in addition the chance of using different storage frequencies depending on the topic and its relevance. This fact not only enhances the granularity of stored message proofs but also provides a more adaptable solution.
Our anti-tampering approach for this accountability proposal comprises two key components: a chain of integrity proofs coupled with the smart contract specification. This section delves into their design and justifies how both align with the constraints of autonomous agents.
Key features such as consistency are guaranteed by the blockchain’s consensus protocol, where the content of each block is hashed and stored in the next one. Additionally, availability is also assured through the distributed and decentralized nature of the blockchain, ensuring accessibility across the entire network.
3.1.1 Chain of integrity proofs
To preserve the integrity of the messages without compromising system performance, our solution entails building a chain where each message hash is linked with the previous hash value. This approach ensures that any alteration of the selected messages will result in a different final hash, easing tampering detection. The verification process is also simplified since only the last hash in the chain needs to be recomputed and compared with the last stored value. A single modification of a hash will disrupt the entire chain from that point onward, simplifying the identification of tampering.
We define a chain entry ($CE$) with a counter $i\in 1..m$ to be linked with the previous one by concatenating the prior digest $h_{i-1}$ with the current log message $Msg_{i}$. hash() denotes the cryptographic hash function, with SHA-256 being the chosen one, commonly used for data integrity verification and digital signatures.
The initial hash value $h_{0}$ of the chain is calculated by applying the selected hash function to a set of $l=32$ random bytes ($RB$), to generate a 256-bit value. Introducing initial entropy into the chain of integrity proofs enhances privacy, ensuring that integrity proofs for similar initial log messages from different records are not identical. Data privacy is further prioritized by storing in the blockchain only the hashes, rather than the contents of the messages.
$CE_{i}$ $=h_{i}$ (1)
$\text{where}~{}RB$$\leftarrow\{0..F\}^{l},$
$h_{0}$ $\leftarrow\text{hash}(RB),$
$h_{i}$ $=\text{hash}(h_{i-1}||Msg_{i})$
By chaining every hash with the previous one, we optimize the validation of stored messages due the last linked digest should correspond to the final proof saved in the blockchain. Additionally, a service has been developed to generate the chain of hashes using the recorded data, ensuring partial validation of the integrity proofs without modifying the stored messages or adding non-essential information to the transactions saved in the blockchain.
The number of proofs included in a blockchain transaction depends on the transaction cost limit set by the blockchain framework. Therefore, being $j\in 1..m$ a transaction counter, each newly arriving chain entry is added in a transaction payload $P_{j}$ until the limit is reached. Being sign() a shorthand for a signature function and $PrivKey$ the private key of the account, transaction $T_{j}$ is transmitted from the service in charge of recording the accountable data to the blockchain.
$T_{j}$ $=(P_{j},S_{j})$ (2)
$\text{where}~{}P_{j}$$\leftarrow\{CE_{i},CE_{i+1},CE_{i+2},\dots\},$
$S_{j}$ $=\text{sign}(P_{j},PrivKey)$
A service interruption or a disruption in the behavior of the autonomous agent will lead to the aggregation of the remaining hash values into a single transaction, even if there is capacity for additional values. This practice guarantees the immutability of the last messages sent by a robot upon completing a task or achieving a goal.
3.1.2 Smart contract
Calling a smart contract function intended for blockchain writing involves the execution of a transaction that changes the state of the storage data on the blockchain. In our approach, this transaction is signed with the account’s private key via the Elliptic Curve Digital Signature Algorithm (ECDSA), effectively identifying the signer.
The defined smart contract also determines its ownership, implementing access control to ensure that only the contract owner can add or update message proofs. These features guarantee authentication and non-repudiation.
Moreover, our smart contract includes a function to verify the existence of a message proof on the blockchain. This fact allows external parties to independently confirm whether a specific message hash has been stored in the contract and when it was stored. This functionality contributes to the system’s transparency, enabling users or auditors to check the validity of message proofs without relying solely on the contract owner, thereby promoting public accountability.
3.2 Explainability Component
Effective communication is essential for achieving explainability in autonomous robots. Our approach is focused on improving this feature through the use of Natural Language Processing (NLP) technology, enabling non-expert users to interact with agents in a meaningful way.
Our solution involves use of RAG to retrieve data from outside the language model (non-parametric) and augment the prompts by adding the information recorded from our accountability component in context. Before this integration, the information undergoes processing by a dedicated module responsible for acquiring and processing the raw data generated by the autonomous agent, serving as contextual input in the question-answering system.
Given the volume of messages, their varied nature, their length, the presence of non-meaningful records generated by the autonomous agent, and the need of establishing specific relationships between messages to enhance the context of the engine, this component extracts relevant information. Its aim is to improve the effectiveness of the explanations provided to non-expert users.
After processing the raw data stored by our accountability component, the ingestion process performs the chunking task to break down extensive text segments into more manageable units. To ensure comparability during relevancy searches at inference time, both this data and the user query are converted into numerical representations through embedding language models. Next, based on the embedding of the user query, its relevant text is identified in the document collection by a similarity search in the embedding space. The user-provided prompt is then extended with relevant information obtained in the search and added to the context. This enhanced prompt is sent to the LLM. As the context contains pertinent external data along with the original prompt, the model output is relevant and accurate.
The addition of new accountable data to our explainability component requires asynchronous updates of the accountable data and its embedding representation. These updates ensure that the model remains sensitive to changes, addressing limitations introduced by offline training.
4 Implementation
The solution presented in Section 3 can be implemented using different technologies. This section details the specific solutions chosen for developing our proposal.
4.1 Black Box Recorder
Robot Operating System (ROS) is the most commonly used framework in robotics, consisting of a set of software libraries and tools for building robot applications. However, challenges in meeting the requirements of real-time embedded systems, and the need to improve its multi-platform capabilities led to a major upgrade to ROS 2, whose main change lies in using the Data Distribution Service (DDS) (Maruyama et al., 2016).
Nodes constitute the fundamental building block in both ROS and ROS 2 applications. They consist of individual processes, each performing specific tasks and communicating with each other via messages. Data can be moved between nodes using different methods, including topics, services, and actions. Topics are used for message passing, while services provide synchronous request-reply interactions. Actions, on the other hand, are needed for long-running tasks that require feedback.
This work focuses on the navigation functionality. To this end, we have used Nav2, a collection of tools for ROS 2, which is considered the successor of the ROS Navigation Stack. Nav2 provides the robot with the ability to perform complex navigation tasks.
Its main functionalities include planning, control, localization, visualization and other additional features. Furthermore, it uses behavior trees to create customized navigation behavior via orchestrating many independent modular servers, which implies high flexibility in the navigation task and the specification of complex robot behaviors (Macenski et al., 2020). Different independent task servers, intended to compute a path, control effort, or recovery, are communicated with the behavior tree (BT) over a ROS 2 interface, such as an action server or service. The use of BTs to generate explanations for robot behavior has been previously successfully evaluated, allowing seeking causal information answer questions (Han et al., 2021).
Previous authors’ work describes the steps involved in the accountable information recording process (Fernández-Becerra et al., 2023). ROS 2 nodes publish data that other nodes consume through ROS 2 topics. Such data include perception information from sensors or commands to actuators from decision-making nodes. Navigation reproducibility of autonomous robots in ROS 2 involves storing essential data, such as their poses, maps, odometry, velocity, and laser scan readings. To this aim, we have used Rosbag files, a binary file format designed to capture and store data generated during the execution of a ROS system. These files can reproduce the exact conditions and inputs from a previous run, making them particularly valuable for understanding system behavior. The set of topics included in the Rosbag files generated are detailed in Table 1.
Table 1: Topics name included in the Rosbag file their description.
Topic name 1 Information provided
map 2-D grid map, in which each cell represents the probability of occupancy.
tf_static Static transform information; fixed relationships between coordinate frames over time.
robot_description URDF (Unified Robot Description Format) or robot model description.
navigate_to_pose/_action/status Information about the current state or progress of the navigation action.
global_costmap/costmap Representation of the environment in terms of costs associated with different areas.
plan Representation of the planned path for the robot to follow as a sequence of waypoints.
rosout Logs and messages from ROS nodes.
local_costmap/costmap Representation of the immediate surroundings of the robot and their cost.
amcl_pose Position and orientation of the robot.
behavior_tree_log Nav2 Behavior Tree nodes status.
cmd_vel Linear and angular velocity commands to the robot’s motion control system.
camera/image_raw Raw image data captured by a camera sensor on the robot.
scan Distances from the sensor to obstacles in the robot’s surroundings.
odom Estimated motion and position of a robot based on sensor data.
tf Coordinate frame transformations in the robotic system.
To ensure the immutability of generated Rosbag files, Ethereum has been selected as the blockchain framework. Ethereum, an opensource public blockchain, introduces the cryptocurrency Ether and a programming platform known as Smart Contract. Employing Proof of Stake (PoS) as a consensus protocol, Ethereum offers advantages in energy efficiency, scalability, security, decentralization, accessibility, and economic efficiency. The high-level algorithms 1 and 2 depict the main steps in recording and proof verification services.
Algorithm 1 Accountable information recording
1:topicsSet = [$selectedTopics$], $serviceCallValue$, $contractAddress$, $contractAbi$, BLOCKS_LIMIT, $privateKey$
2:$rosbag$
3:Boolean $recording\leftarrow$ false
4:String $initialNonce\leftarrow$ sha256($randomBytes$)
5:// Recording process attending to a service call.
6:$recording\leftarrow serviceCallValue$
7:if $recording$ = $true$ then
8:     String $previousHash\leftarrow$$initialNonce$
9:     for $message$ in $topicsSet$ do
10:         if $messageNumber$ $topicRate$ == 0  then
11:              $chainedHash\leftarrow$ sha256($message$,$previousHash$)
12:              $previousHash\leftarrow$ $chainedHash$
13:         end if
14:         // Building chain of hashes.
15:         if $payloadSize$ < BLOCKS_LIMIT then
16:              $payload\leftarrow$ $payload\cup chainedHash$
17:         else
18:              // Load Solitidy contract.
19:              loadBcContract($contractAddress$, $contractAbi$)
20:              // Saving chain of hashes into blockchain.
21:              $transaction\leftarrow$buildTransaction($payload$)
22:              $signedTransaction\leftarrow$ signTransaction($transaction$, $privateKey$)
23:              sendTransaction($signed\_transaction$)
24:              for $hash$ in $payload$ do
[…]
26:              end for
27:         end if
28:     end for
29:end if
30:return $rosbag$
Algorithm 2 Proof verification
1:$hashValue$
2:$BlockNumber$
3:// Load Solitidy contract.
4:loadBcContract($contractAddress$, $contractAbi$)
5:$transaction\leftarrow$buildTransaction($hashValue$)
6:$signedTransaction\leftarrow$ signTransaction($transaction$, $privateKey$)
7:sendTransaction($signed\_transaction$)
8:$blockNumber\leftarrow$readBC($hashValue$)
9:if $blockNumber$<>0 then
10:     $msg\leftarrow$ “The hash value is stored in $blockNumber$."
11:else
12:     $msg\leftarrow$ “The hash value is not stored."
13:end if
14:return $msg$
Our approach addresses challenges such as the drop in robot performance during the Rosbag recording process and costs associated with storing blockchain data. We accomplish this by saving chained hashes from messages at different intervals. These intervals can be established based on the relevance of messages from a topic, regarding their required immutability features. Furthermore, message proofs are stored in real-time on the Ethereum blockchain, eliminating the need to modify the content of the Rosbag file. Detection of any alteration to the messages within this file is facilitated by a service that generates the hash chain from the file at intervals set for each topic during the recording process. To verify the existence of any of these hashes in the blockchain, the blockchain logic has been extended by developing a Smart Contract (SC) using the Solidity language. Solidity is an open-source programming language that facilitates the creation of decentralized applications, contracts, protocols, and agreements. This contract provides functionalities for storing and verifying proof information for hashes stored on the Ethereum blockchain by associating each hash value with the number of block where it is stored. Verification of transaction ownership is also conducted during the storage of these proofs.
Therefore, this approach combines immutability, decentralization, traceability, and cost efficiency to address challenges related to the real-time recording and storage of Rosbag files securely and transparently.
4.2 Log Curation Component
To filter relevant information from the Rosbag file messages, a raw data processing node has been implemented. Given the size, variety, and intricate nature of messages contained in the Rosbag file, the inclusion of this raw information might be challenging for non-expert users. Therefore, it is necessary to include a component that processes this data, eliminating non-essential content in the natural language explanation generation process. By analyzing messages from the recorded Rosbag topics, this service can identify key circumstances in the navigation process, such as the appearance of an obstacle or a change in the planned trajectory. As a result, through this component, ROS 2 developers are relieved from having to call the logging API to generate explanations. This functionality streamlines the development process, allowing developers to focus on core functionalities without need for managing log calls. This approach reduces debugging efforts and boosts efficient resource usage, scalability, and adaptability in the development lifecycle, contributing to a more productive natural language explanation generation process. An example of how raw ROS messages are transformed through this component is included in Figure 2.
Rosbag Raw Messages action_msgs.msg.GoalStatusArray(status_list=action_msgs.msg.GoalStatus(goal_info= action_msgs.msg.GoalInfo( goal_id=unique_identifier_msgs.msg.UUID(uuid=array([190, 81, 48, 208, 248, 147, 78, 82, 188, 116, 202, 152, 203, 225, 149, 158], dtype=uint8)), stamp=builtin_interfaces.msg.Time(sec=27, nanosec=933000000)), status=2)]) nav2_msgs.msg.BehaviorTreeStatusChange(timestamp=builtin_interfaces.msg.Time(sec=1701354686, nanosec=130104968), node_name=’ComputePathToPose’, previous_status=’IDLE’, current_status=’RUNNING’) Processed Data 1701354686 Navigation to the goal number 1 has started. 1701354686 Navigation to the goal number 1 is in progress. 1701354686 Nav2 Behavior Tree node NavigateRecovery that recovers from unexpected situations, is running.
Figure 2: Example of raw Rosbag messages and information obtained from them after the Log Curation process.
4.3 Explainability Component
The output of the above described node serves as an external source in our implementation of the RAG approach. This development is based on the use and extension of localGPT  [^1] https://github.com/PromtEngineer/localGPT, an open-source question-answering system designed to interact with documents conversationally while safeguarding the privacy of the underlying information. Running locally, it ensures data privacy and secure interactions with the interpreted and recorded data, facilitating the generation of natural language explanations. Its functionality involves loading a set of selected documents, breaking them into smaller chunks, augmenting search capabilities with LangChain tools, and generating local embeddings. The outcomes are stored in a local vector database using Chroma vector storage. Once this information is saved, and given a user input, relevant splits are retrieved from storage via a similarity search, which finds the appropriate piece of information from the documents. The retrieved data is used by the selected LLM when producing an answer. Figure 3 depicts the main components in the RAG architecture used.
Figure 3: Retrieval Augmented Generation arquitecture.
Therefore, this implementation combines the strengths of robotics frameworks, blockchain technology, and natural language processing to provide a comprehensive solution for real-time recording, secure storage, and user-friendly interpretation of robotic system data. It addresses challenges related to performance, storage costs, and user comprehension, making it an effective approach when understanding the behavior of autonomous robots.
4.4 Software and hardware artifacts
Our solution has been deployed in a Linux server running Ubuntu 22.04 as Operating System (OS). The server is equipped with 6-core 2.6 GHz CPU, 64 GB of memory, and a 4 GB GPU.
The chosen ROS 2 version is Humble. Details regarding the topics incorporated into the Rosbag file, along with their average rates, are provided in Table 2. These topics play a crucial role in ensuring the reproducibility of the navigation process, facilitating data analysis and debugging purposes.
The experimentation was executed entirely in Gazebo, an open-source solution for 3D robotics simulations. To provide a realistic scenario, we used a hospital simulation environment made available by Amazon Web Services (AWS) [^2] https://github.com/aws-robotics/aws-robomaker-hospital-world. Additionally, we selected RB-1, an autonomous mobile robot based on the ROS 2 platform produced by Robotnik (Guzmán et al., 2017). This combination of tools and platforms enabled us to assess the impact and effectiveness of our approach.
The interaction with Ethereum has been developed through web3.py library (version 6.6.1). Ganache (version 2.7.1) has been used as the Ethereum network simulator.
RAG solution holds Langchain (version 0.325), ChromaDB (version 0.4.6). We used the open source model Llama 2 7B GGUF (Touvron et al., 2023) as the LLM responsible for generating natural language explanations for end users. This LLM integrates RLHF and shows versatility in handling general-purpose tasks, ranging from answering questions, text generation, and language translation to more demanding tasks such as reasoning or coding. Preliminary tests of Llama 2, compared to other LLMs such as Mistral (Jiang et al., 2023) or Zephyr (Tunstall et al., 2023), led us to choose this model for evaluating our data, due to the promising results obtained in these assessments conducted during the last quarter of 2023, when this work was in its development phase. As text embedding model we have chosen Instructor-xl (Su et al., 2022), a high-performance model designed to generate text embeddings tailored to tasks such as classification, clustering, text evaluation or retrieval, fact particularly important in the processes performed for any RAG solution.
Table 2: Robag topics rate and average rate per message
Topic Avg. Rate (Hz) Avg. Size Msg. (B)
map 0 1007308.8
tf_static 0 9011.2
robot_description 0 25497.6
navigate_to_pose/_action/status 0.028 2389.333
global_costmapcostmap 0.500 996147.2
plan 0.884 3809.28
rosout 1.014 746.057
local_costmapcostmap 1.646 3986.99
amcl_pose 1.925 1054.72
behavior_tree_log 2.74 313.224
cmd_vel 18.151 147.110
camera/image_raw 25.57 1037110.93
scan 47.173 8846.012
odom 92.946 779.618
tf 387.278 195.413
5 Evaluation
To evaluate our proposal, we conducted an experimental study comparing the performance of an accountability system built in accordance with Section 4. This study covered several approaches for storing integrity proofs across various scenarios. Additionally, we explored and adapted the capabilities of RAG solutions and LLMs in question-answering systems, using as source the data gathered by our black box-like component, through a comprehensive set of questions. This work aimed to asses the effectiveness of these technologies in delivering natural language explanations to individuals who are not experts.
The experiments conducted within this proposal focus on navigation tasks included in the Navigation Functionality of the European Robotics League Consumer Service Robots Challenge (ERL Consumer) (ERL, 2023).
This functionality evaluates the capabilities of a robot to navigate accurately and autonomously within a given scenario, containing furniture, objects, and people. Details like the initial position and the number of waypoints the robot needs to visit are provided beforehand. Furthermore, the coordinates of the waypoints are communicated sequentially to the robot during runtime. Each waypoint is characterized by its $X$, $Y$, and $\Theta$ coordinates. Existing elements in the environment, such as furniture, doors, or walls, must be mapped. The obstacles, with varying shapes and sizes, remain unknown to the teams beforehand and may differ between runs.
Our experiments are structured around three distinct assessment scenarios in line with the navigation functionality previously described. In the initial scenario, the robot must navigate to three waypoints, encountering no obstacles along the way. The second scenario presents the robot with the challenge of adjusting its pre-planned path to reach the second goal. This adjustment becomes necessary due to the appearance of an obstacle obstructing the original planned route, requiring a deviation to achieve the goal. The third scenario introduces a new obstacle that blocks the only door needed to reach the first goal. Consequently, this obstacle prompts the cancellation of the first waypoint objective. The second and third goals are achieved with the same restrictions as detailed in the second scenario.
The three scenarios are represented in Figure 4.
The entire source code is accessible online on GitHub [^3] https://github.com/laurafbec/immutable_explainable_BBR.git, and is licensed under GPLv3.
Figure 4: Scenario maps with trajectories, goals, and obstacles. (a) Scenario 1: No obstacles scenario. (b) Scenario 2: One obstacle scenario. (c) Scenario 3: Two obstacles scenario.
5.1 Black Box Recorder evaluation
In evaluating the accountability component, we conducted six experiments across three previously described scenarios. These experiments involved integrity proofs storage in the blockchain, with hash calculations occurring at intervals of one every 10, 25, 50, or 100 messages. Moreover, our solution introduces the flexibility to save proofs of evidence at different rates based on the topic, ensuring non-tampering for critical topics or those with lower publishing rates where message loss could be crucial.
Therefore, our fifth recording method is designed to store a hash for every message within a topic with a rate below 0.5 Hz, one hash for every 5 messages for topics with rates below 1 Hz, a hash for every 10 messages if the rate is below 3 Hz, one hash for every 15 messages if the topic’s rate is below 25 Hz, a hash for every 50 messages for topics with rates below 45 Hz, and one hash for every 100 and 1000 messages when the rate falls below 100 Hz and 400 Hz, respectively. This approach, together with those operating at fixed intervals, has been compared with the generation of a Rosbag file that covers identical topics but lacks anti-tampering properties. Each experiment has been repeated six times, totalling 108 runs, to derive insights into the characteristics of the black box-like device.
5.1.1 Metrics
We employed different metrics to evaluate the impact of the accountability system in the chosen scenarios. First, through the topics’ publication rate and the number of messages stored per topic in each Rosbag file, we analyzed the average rate of lost messages based on the employed blockchain storage approach.
Next, regarding system performance, we studied variables such as CPU, memory and disk usage. Specifically, we examined the following variables: system load 1-, 5- and 15-minute average, CPU usage (%), memory usage (GB), sent network traffic (MB/s) and disk writing rate (MB/s). These variables are highly significant for system-overhead analysis given that elevated load conditions can potentially influence the robot’s behavior and event-recording capabilities. In particular, we have obtained a full report about system load, CPU, RAM, network traffic and disk write operations updated every second, facilitating the tracking of overall system activity and an easy comparison with preceding values. To describe the overall distribution and characteristics of these results, we calculated their average as statistical measure. This information is included in Section 6.
5.2 Explainability evaluation
Despite the remarkable performance of LLMs across diverse language tasks, evaluating these models presents challenges, especially in ensuring factual consistency. The extensive capabilities of LLMs and the limitations of current benchmarks in capturing human preferences, have given rise to emerging trends in LLM evaluations, such as employing strong LLMs as judges (Zheng et al., 2023). Traditional similarity-based metrics, such as BLEU and ROUGE, have proven to misalign with human evaluation. While human evaluation remains the primary standard for assessing human preferences, it is exceptionally time-consuming and expensive. To automate the evaluation process, several works have adopted the ’LLM-as-a-judge’ approach, which focuses on using strong language models (LLMs) like GPT-4. These models are often trained with reinforcement learning from human feedback (RLHF), resulting in strong human alignment (Lin, 2024). Previous studies have evidenced the consistent agreement of these models with human grading in over 80% of judgments (Bubeck et al., 2023; Helgesson Hallström, 2023; Leng et al., 2023). Given these characteristics, we opted for GPT-4 to assess the answers generated by our explainability component. Through LLM-as-a-judge, we are able to obtain not only scores but also explanations, making their outputs interpretable. The whole process is represented in Figure 5.
Figure 5: Explainability evaluation.
LLMs exhibit a certain degree of indeterminism, defined as the model’s ability to offer different answers when asked the same question or prompt several times. This behavior derived from their training, in which the models learn to generate outputs by sampling from probabilistic distributions over possible word sequences, introducing randomness to their response. The effect of this indeterminism must be considered in any LLM evaluation.
In addition, RAG for question-answering systems is one of the most difficult LLM use cases to evaluate. A basic RAG setup includes two main sources of error: the retriever and the LLM. The evaluation of the retriever aims at determining the relevance of its document selection in relation to a query. To this end, we conducted an experimental study running with 16 different questions over the data generated and processed during the navigation task in the three previously described scenarios. This set of questions has been evaluated six times in each scenario, giving a total of 288 questions and answers to perform the analysis. This approach ensures a thorough examination of the performance of the RAG system in various scenarios, showing potential strengths and weaknesses in our solution.
Using the context provided by the Rosbag data and an evaluation question, our approach includes an evaluation component for the generated explanations that contrasts the prediction made by the evaluated LLM and, thanks to the retrieved selection, compares it to an annotated ground truth. This process yields a correctness metric that scores whether the retriever returned no relevant documents (0) or whether it returned a perfect selection of documents (1). The final score, an average of all the scores in the dataset, is used to judge the retriever’s overall performance.
A comprehensive evaluation of a LLM should consider multiple metrics that include correctness as well as other aspects. A language model should produce text that is not only correct but also logically structured and easy to understand. Produced response must be appropriately detailed without unnecessary verbosity. Furthermore, the LLM must not only provide correct information but also delivers in a meaningful and valuable way to the user.
Therefore, to assess the performance of the selected LLM on specific attributes, our approach also includes a criteria-based evaluation covering accuracy, conciseness, coherence, and helpfulness. Selected criteria metrics are describe in Table 3. This structured assessment offers a means to assess the model’s effectiveness in generating responses, providing a broader understanding of the model’s capabilities and its alignment with the used scenarios. Employing a scorecard defined for each of these features, our solution assigns scores between 0 and 10 to the responses it generates. The use of this scorecard facilitates a detailed and quantitative analysis of the LLM’s output, ensuring a comprehensive understanding of both its strengths and areas for improvement. A complete example of the answers and evaluations obtained for a selected question in the three selected scenarios is included in Figure 6.
Table 3: Criteria-based evaluation metrics
Criteria metric
Description
Accuracy
Accuracy evaluates the extent to which the response conforms to the reference, ranging from being completely unrelated to being completely accurate and in line with the reference.
Helpfulness
Helpfulness evaluates the degree to which the response provides valuable information, ranging from providing no helpful information to being extremely helpful and offering comprehensive information.
Conciseness
Conciseness assesses the brevity of the response, ranging from being extremely verbose to highly concise with no unnecessary words.
Coherence
Coherence examines the organizational structure and logical flow of the response, from extremely inconsistent to highly coherent, with a clear and logical organization.
Question: How many goals have been reached by the robot? Scenario 1 Answer: Based on the provided context, we know that the robot has reached 3 goals. The context mentions that the goal number 1, 2 and 3 have been reached successfully. Therefore, the answer to your question is 3. GPT-4 Judgement Predicted Grade: CORRECT Reasoning: The AI assistant’s response is accurate, concise, coherent, and helpful. It directly answers the user’s question by confirming that the robot has reached three goals. The assistant also provides clarification by referencing the context, which adds to the response’s accuracy and helpfulness. There are no unnecessary words or unclear statements, making the response concise and coherent. Accuracy: [[10]] Conciseness: [[10]] Coherence: [[10]] Helpfulness: [[10]]", score: 1.0 Scenario 2 Answer: Based on the context provided, the robot has reached 3 goals. GPT-4 Judgement Predicted Grade: CORRECT Reasoning: The assistant’s response is completely accurate as it matches the information given in the ground truth. The response is concise, with no unnecessary words or information. It is highly coherent and has a clear structure. The response is extremely helpful as it gives the exact information asked for by the user. Therefore, the ratings are as follows: Accuracy: [[10]] Conciseness: [[10]] Coherence: [[10]] Helpfulness: [[10]]", score: 1.0 Scenario 3 Answer: Based on the provided context,the robot has successfully reached 2 out of 3 goals. The successful goals are: * Goal number 2: Navigation to the goal number 2 has succeeded. Position: -9.14068582195674, -25.977117025927537. Orientation: -0.6580310307304338,0.7529908117605705. * Goal number 3: Navigation to the goal number 3 has succeeded. Position: -1.8811226231679194, -28.773672979732826. Orientation: -0.5946646510401259,0.803973850820613. The aborted goal is: * Goal number 1: Navigation to the goal number 1 has been aborted. Position: -5.062160931312302, -8.330878461767831. Orientation: -0.8760794639388502,0.4821667479872674. Therefore, the answer to the user’s question is: The robot has successfully reached 2 out of 3 goals. Goal numbers 2 and 3 have been successfully achieved, while goal number 1 has been aborted. GPT-4 Judgement Predicted Grade: CORRECT Reasoning: The AI assistant’s response is comprehensive and accurate. It correctly states that the robot has successfully reached goals number 2 and 3 while goal number 1 has been aborted. The assistant also provides additional details about the positions and orientations of the goals. While these extra details are not explicitly asked for in the user’s question, they may be useful for the user’s understanding. Therefore, the assistant’s response is also helpful. However, the response is not concise because it includes more information than necessary, leading to some redundancy in the explanation. The assistant’s response is coherent as it logically organizes the information about each goal. Accuracy: [[10]] Conciseness: [[7]] Coherence: [[10]] Helpfulness: [[10]]", score: 1.0
Figure 6: Examples of correct answers, RAG, and criteria-based evaluations obtained across the three scenarios for the question ’How many goals have been reached by the robot?’
5.2.1 Categories and evaluation questions
To evaluate our explainability proposal in the context of ROS 2 messages interpretation, we have defined a set of questions covering diverse aspects of the navigation process over the how, the what, the which, the when and the where dimensions. These questions provide insights into different stages and components of the three scenarios. To structure and improve the understanding and analysis of this set of questions, we have organized them into five different categories. These categories range from high-level overviews of navigation to detailed inquiries about trajectory planning, goal completion, and the functioning of the behavior tree. The categories also delve into the causes of unexpected behavior or errors in navigation, improving the system explainability whenever an unexpected action occurs. Through these questions, we aim to address key concerns in XAI, including understanding why decisions are made, establishing causality between actions, determining success or failure, identifying timing, and disguising the methods by which achievements are realized (Gunning and Aha, 2019).
The set of categories, their descriptions and the questions included in each one are described next.
Navigation Process Overview
These questions provide an overview of the navigation process captured in the Rosbag file. Understanding what has happened and how the navigation task has proceeded provides a high-level context that sets the stage for more detailed questions in later categories. This category comprises the following questions:
  1. 1.
    What has happened in this ROS2 log regarding navigation?
  2. 2.
    How has the navigation task proceeded?
Trajectory Planning and Replanning
These questions delve into the specifics of trajectory planning, a critical aspect of navigation. They address whether the robot has re-planned its trajectory, the reasons behind this behavior, and whether it has encountered any obstacles. These details are necessary for understanding how the robot dynamically adjusts its path. The questions included in this category are:
  1. 3.
    Has the robot re-planned an alternative trajectory during navigation?
  2. 4.
    Why did the robot re-plan the route?
  3. 5.
    Did the robot find any obstacle during the navigation?
Goal Completion and Navigation Task Status
These questions focus on the completion status of the navigation goals. Knowing how many targets have been reached, whether the robot has completed the navigation task, and when the navigation task ended provides insights into the overall progress and success of the robot’s task. This category includes the following questions:
  1. 6.
    How many goals have been reached by the robot?
  2. 7.
    Has the robot completed the navigation task?
  3. 8.
    When has the robot ended the navigation task?
  4. 9.
    Have all objectives been successfully achieved or have any been cancelled or aborted?
Specifics about Goals and Locations
These questions examine the specifics of goals and locations. Understanding the location of goals, the robot’s linear velocity during navigation, its initial position and orientation, and the velocity after receiving a goal contribute to a detailed understanding of the robot’s movements and performance. Questions included in this category are:
  1. 10.
    Where is the second location or goal located?
  2. 11.
    What was the linear velocity when navigating to goal pose number 2?
  3. 12.
    What is the initial position and orientation of the robot?
  4. 13.
    What was the linear velocity of the robot after receiving the goal number 1?
Nav2 Behavior Tree and Node Status
These questions provide information about the underlying behavior tree and the state of its nodes in the navigation system. Knowing the specific node responsible for determining a viable path, identifying any failed nodes during navigation, and understanding the use of behavior tree nodes, contribute to a deeper understanding of the inner workings of the navigation system. This category includes:
  1. 14.
    What is Nav2 Behavior Tree’s node to determine a viable path from a starting point to a specified target pose or location?
  2. 15.
    Did any node from the Nav2 Behavior Tree fail during navigation?
  3. 16.
    Which Behavior Tree’s nodes were used during navigation?
6 Experimental results and discussion
This section summarizes the experiments’ results involving both the black box-like engine and the explainability component when using our solution in the previously depicted assessment scenarios.
6.1 Message loss rates
Figure 7 shows the percentage of lost messages in the Rosbag file for experiments detailed in Section 5, capturing the topics outlined in Table 2, attending to their publication rates and to the approaches selected to store integrity proofs in blockchain. The results across the three scenarios demonstrate a consistent trend, showing uniform message loss patterns in the creation of immutable Rosbag files.
Choosing to store a single message hash for every 100 messages results in a loss of less than 5% compared to avoiding anti-tampering measures in the Rosbag recording process, even for topics with rates exceeding 300 Hz. This approach stands out as the most efficient option for integrating anti-tampering techniques into our accountability solution.
Alternative strategies, such as saving one proof every 50 messages or adjusting proof intervals based on topic rates, achieve losses below 10% for each topic. This is particularly noteworthy, ensuring the integrity of messages based on their significance in a given scenario or task.
Introducing an integrity proof every 25 messages leads to an overall reduction in message recording, especially evident when topic rates surpass 100 Hz. Storing one hash every 10 messages could lead to losses exceeding 30% for topics with rates higher or around 100 Hz. Beyond the consistent linear trend of message loss with high topic rates, a slight deviation is observed in this approach when topics fall below 50 Hz. This variation signifies the influence on message recording when computing hash functions for messages nearing 1 MB in size, as observed in messages from the /camera/image_raw topic, given that the computational complexity of common hash functions, such as SHA-256, increases with input size.
Figure 7: Message loss rates in Rosbag Files under several anti-tampering techniques. Comparative visualizations for: (a) Scenario 1, (b) Scenario 2, and (c) Scenario 3.
6.2 System benchmarking
The system load or, the running thread demand on the system as the average number of running and idle threads, is shown in Figure 8. The average of 1 minute is higher than the average of 5 or 15 minutes in all scenarios, showing a load increase. However, the values of this metric do not exceed the CPU count, which is 6 in our experimental setup, indicating no performance issues or resource saturation. The inclusion of runnable and uninterruptible tasks in this metric means that the load average may increase due to a disk I/O workload, not just CPU demand, giving a more general view of system overhead. Other more specific metrics are described next to clarify the possible existence of a bottleneck in our approach.
Figure 8: Load average under several anti-tampering techniques. Comparative visualizations for: (a) Scenario 1, (b) Scenario 2, and (c) Scenario 3.
The CPU consumption by user processes maintains a consistent linear trend across all approaches and scenarios, remaining around 30%. This level of consumption is generally considered acceptable for resource-intensive processes such as the recording of a Rosbag file, providing a substantial margin in the event of introducing other critical tasks into the system.
Concerning memory consumption, all approaches within the three scenarios exhibit comparable values, differing by no more than 0.5 GB. Consequently, the inclusion of blockchain approaches into our black box recorder does not yield a remarkable impact on RAM comsuption in comparison to avoiding the use of anti-tampering techniques in this recording process.
Regarding network traffic, approaches with higher process storage in the blockchain increase this metric, but do not exceed 0,20 MB/s, which is a very affordable quantity. Furthermore, approaches such as storing one hash every 100 messages nearly match the avoidance of storing any proof into the blockchain, indicating that including this approach do not have a remarkable impact in the system performance.
Disk write operations do not form a bottleneck in system performance, with all approaches differing by no more than 0,40 MB/s, which is not a very significant impact. In addition, all the approaches that perform integrity proof storage on the blockchain achieve lower or similar results to the approach avoiding anti-tampering operations.
Figure 9 shows the results described.
Figure 9: System benchmarking results: (a) CPU usage of user processes. (b) RAM usage. (c) Sent network traffic. (d) Disk writes traffic.
These findings may encourage the adoption of our approach to ensure integrity in messages from autonomous agents.
6.3 Correctness evaluation
Figure 10 shows the results of the correctness of RAG evaluation, presenting both the overall results and assessments based on predefined categories. The results demonstrate remarkable consistency across the three scenarios, getting a significant number of correct answers, even though there has not been a prior tuning in the ROS 2 messages for autonomous agents navigation tasks. The slight percentage of incorrect answers remains within an acceptable threshold, given the specific context employed and the inclusion of general questions.
Figure 10: (a) Overall correctness in the three scenarios. (b) Correctness by category in the three scenarios.
Regarding individual categories, the correctness in the category ’Navigation Process Overview’ is relatively consistent across all scenarios, indicating a good understanding of events and progress related to navigation. However, the broad nature of the question "What has happened in this ROS2 log regarding navigation?" initially yielded a correctness below 65% in the third scenario. In this instance, the cancellation of a goal resulted in the execution and failures of Nav2 Behavior Tree nodes, not included in the other two scenarios and focused on navigation recovery. Therefore, the incorporation in a dataset of information related to the cancellation of a goal, results in LLM asnwers that usually diverge from the ground truth provided, as the model occasionally prioritized Behavior Trees information, other times focused on locations or timestamps. This variability made it difficult to align with the ground truth provided, even if the obtained answers could be considered partially accurate. Introducing the prompt extension "Describe the sequence of events regarding navigation goals and include information about each of them." in question 1 significantly increased correctness in the first category, achieving results around and exceeding the 75% in all three scenarios. This fact underscores the effectiveness of prompt engineering when tackling ambiguous or broad queries.
The results also show consistently high correctness across all scenarios concerning questions related to trajectory planning and re-planning. The model is able to determine whether the modification of a predicted trajectory has been caused by the occurrence of an obstacle or other circumstance. Scenario 1 and Scenario 2 also exhibit relatively high correctness in questions related to goal completion and task status. However, the model encounters challenges in Scenario 3, occasionally failing to identify if a navigation that starts correctly concludes with an abortion or cancellation. This issue may be attributed to misalignment during the retrieval process, specifically in the selection of the best chunk associated with a particular goal associated with numerous embeddings within a single collection from the vector database.
Correctness in identifying specific targets and locations lacks some consistency, with results exceeding 50% in the first scenario, around 70% in the second and over 75% in the third. In particular, three scenarios encounter difficulties in identifying the initial position of the robot before starting the navigation, due to the model misaligns the queried position with the one assumed by the robot when starting the navigation process. This fact may also be attributed to deficiencies in the retrieval process. In Scenario 1, understanding the sequence of events or the order of goals results challenging for the model, leading to inaccuracies in specifying the second location or goal. Nevertheless, this challenge is not present in Scenario 3 and absent or reduced in Scenario 2, demonstrating how differences in data distribution can introduce varied responses even when using specific and unambiguous questions.
Through questions included in the Nav2 Behavior Tree and node status category, we assessed the effectiveness of the model’s formal linguistic competence compared to its functional linguistic competence, a feature of present LLMs (Mahowald et al., 2023). While the model excels in identifying Behavior Tree Nodes given its description, it faces challenges in pinpointing these components under specific conditions. As a result, the category correctness suffers a drop in Scenario 3, although overall results remain above 75%. In this scenario, the navigation recovery actions caused by the cancellation of the first goal introduced a challenge in the interpretation of the information by the LLM due to the execution of several Nav2 Behavior Tree nodes not present in the others scenarios. Particularly, the mistakes stem from the model’s confusion in determining whether a previously failed Nav2 Behavior Tree node has recovered or finished its execution, as well as identifying all nodes that were executed during navigation. This can be attributed to the fact that the complexities of the recovery actions and the dynamic nature of Behavior Tree executions during navigation recovery introduce complexities that the model may not fully generalize from the training data, underscoring the need for additional fine-tuning in such scenario-specific contexts in case the obtained correctness needs to be improved.
6.4 Criteria-based evaluation
Figure 11 shows the remarkable performance of the Large Language Model’s (LLM) in the evaluated metrics, demonstrating consistency with slight variations in the scores. On average, the answers provided are accurate and in line with the reference, though there might be room for improvement. While responses generally provide useful information, there are instances where additional depth could enhance their utility. In addition, most answers usually include a remarkable balance between conciseness and effectiveness, showing a proficient level of brevity, as well as clear and logical organization. The overall results show no notable differences in these metrics across the three scenarios. This suggests that messages within the Rosbag related to the introduction of obstacles or cancellation of goals do not notably influence the general criteria evaluation outcomes.
Figure 11: (a) Overall criteria evaluation in the three scenarios.
The language model’s adaptability to challenges like obstacles and goal cancellations regarding the described categories, is shown in Figure 12.
Figure 12: Large Language Model criteria-based evaluation by categories. Comparative visualizations for: (a) Scenario 1, (b) Scenario 2, and (c) Scenario 3.
Concerning general questions about about events and the progression of the navigation task, we observe variations in scores across Scenarios 2 and 3. While coherence and conciseness are even improved, Scenario 1, involving direct navigation to three goals, outperforms in accuracy and helpfulness. This fact suggests that effective communication on general queries benefits from the absence of obstacles or goal cancellations, enhancing overall understanding.
In the category ’Trajectory Planning and Replanning’, the questions are effectively addressed in all the three scenarios, showing scores consistently above 8 in every evaluated metric, even when obstacles are introduced or a goal is cancelled. These results underlines the system’s proficiency in handling trajectory-related questions despite potential new challenges.
Regarding the category ’Goal Completion and Navigation Task Status’, Scenarios 1 and 2 demonstrate the effectiveness of the system in conveying goal completion and navigation task status information. Nevertheless, the slightly lower scores for Scenario 3 indicate potential challenges in effectively communicating completion status when a goal is aborted.
Concerning the ’Specifics about Goals and Locations’ category, the system consistently provide valuable goal and location information, even in the presence of obstacles or goal cancellations. This suggests a robust capacity to understand and localize specific details regardless of challenges introduced in each scenario.
The results regarding questions included in ’Nav2 Behavior Tree and Node Status’ demonstrate effectiveness in communication across all scenarios and metrics. This emphasizes the potential for Behavior Trees to complement language models in explainability, offering insight into the decision-making processes during navigation tasks.
In summary, although results can be improved in accuracy regarding general questions, the language model exhibits adaptability to different navigation scenarios, facing obstacles and changes in navigation tasks. The consistency in conveying information related to Nav2 Behavior Trees suggests a robust understanding of underlying navigation processes. The model also performs well when providing information about trajectory planning, goal completion, and specifics about goals and locations, regardless of the specific challenges introduced in each scenario. Its robustness in effective communication, despite technical vocabulary, details and the introduced challenges, is demonstrated through the achievement of responses that are not only generally accurate but also easily comprehensible to users without expertise, which enhances the model’s practical utility.
7 Conclusions
This work presents an accountability and explicability architecture, implemented for ROS-based mobile robots. Our approach comprises a black box component for accountability that effectively addresses real-time recording challenges, ensuring message integrity and verifiability through blockchain technology, and an explainability component. This last component is based on the use of RAG and LLM through a private question-answering system designed to use the information recorded by the first component as context. This solution demonstrates the feasibility of obtaining accurate, coherent, concise and helpful explanations for robot actions from raw data stored in Rosbag files, thereby enhancing human-robot interaction. Additionally, both main components have been integrated through the development of a third main functionality in charge of filtering non-relevant information and making messages closer to the final user, due to their specific nature can introduce challenges in their comprehension process. Furthermore, this last component releases software developers from making direct calls to the logging API, easing the development process and improving the scalability and adaptability of our solution for every navigation task performed by a ROS-based robot.
The results achieved highlight the capability of our solution to guarantee the immutability of information generated by an agent without compromising its performance, a feature particularly important when trying to identify the causes behind an event. Moreover, our system improves efficiency and effectiveness in the deployment of autonomous agents, providing natural language explanations across key navigation aspects including goal achievement, obstacle appearance, trajectory re-planning, and objective cancellation or abortion. This fact strengthens trust and confidence in the use of autonomous agents across diverse applications and contexts.
While our approach has shown remarkable results in navigation functionalities regarding several scenarios and circumstances, future work should involve adapting our solution to other applications by simply adapting the recording and the processing components to the relevant topics of the area or domain of use. Furthermore, the consistency and robustness of responses generated by our explainability approach could benefit from refinement across different scenarios, especially when categorizing questions under specific subjects. Therefore, future work should explore the impact of In-Context Learning (ICL) techniques on prompt generation, in addition to fine-tuning the model for the specific context of autonomous agent navigation tasks.
Additionally, our RAG approach could improve in terms of retrieving the most relevant chunk of information from the context. This enhancement may involve exploring the use of separate collections within the vector database or employing a hybrid approach, such as integrating knowledge graphs (KG) with vector indexes. These approaches can help to improve the accuracy, quality and overall consistency of the answers obtained across several scenarios, thereby enhancing the performance of our proposed system.
[…]
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Data availability
Data and tools are available in a public repository.
Acknowledgments
This research has been partially funded by the Recovery, Transformation, and Resilience Plan, financed by the European Union (Next Generation) thanks to the TESCAC project (Traceability and Explainability in Autonomous Systems for improved Cybersecurity) granted by INCIBE to the University of León; and by grant PID2021-126592OB-C21 funded by MCIN/AEI/10.13039/501100011033 EDMAR (Explainable Decision Making in Autonomous Robots) project, PID2021-126592OB-C21 funded by
MCIN/AEI/10.13039/501100011033 and by ERDF ”A way of making Europe”.